# Band-dominated and Fourier-band-dominated operators on locally compact abelian groups

### v1

@@ Introduction
where $e$ is the neutral element of the group $\Xi$. For $G:=\mathbb{R}^{d}$ (and hence $\Xi=\mathbb{R}^{2d}$), $\mathcal{C}_{1}(\mathcal{H})$ has spectral properties similar to band-dominated operators, cf. [9, 11]. In fact, it is well known that in case $G=\mathbb{Z}^{N}$ (hence $\Xi=\mathbb{Z}^{N}\times\mathbb{T}^{N}$) the algebra of band-dominated operators actually agrees with $\mathcal{C}_{1}(\ell^{2}(\mathbb{Z}))$, cf. [21, Theorem 2.1.6].
@@
Indeed, what we are going to prove is a bit more detailed than the above statements, and the precise formulation needs some more technicalities that we only mention in later parts of the paper. To briefly summarize this, we prove the result not only for the Hilbert space $L^{2}(G)$, but also for operators on the modulation spaces $M^{p}(G)$. If the group $G$ is discrete, then $M^{p}(G)=\ell^{p}(G)$, putting the results in alignment with the classical results for band-dominated operators over sequence spaces, noting that $\operatorname{BDO}(M^{p}(\widehat{G}))=\mathcal{L}(M^{p}(\widehat{G}))$ and $\partial\Xi=\partial G\times\widehat{G}$ in this case. When the group is not discrete, then $M^{p}(G)\neq L^{p}(G)$ for $p\neq 2$. Further, we are actually able to classify the band-dominated operators (and not only their intersection with Fourier-BDO) in terms of continuity with respect to a certain group action, leading to the above theorem. Let us also mention that we do not assume that $G$ is metrizable. In fact, we do not even need first countable. This contrasts to previous work such as [14] and [23], where the metric is an essential part of the theory. Instead of a metric, we will use the uniform structure that the lca group provides. We note that this involves some technical issues when the group is not second countable. Nevertheless, all these problems can be overcome, mainly by making an appropriate change in the definition of what the space $L^{\infty}$ is. We shall not elaborate on the details of this here and instead refer to the discussions in [10, End of Section 3] or [7, Section 2.3].
The paper is organized as follows. In Section 2, we describe the basic setting of the present work, combined with an introduction to coorbit spaces, as suitable for the precise setting we are working with. Since we also want to make this section available as a reference for later work, we put some more details there than would be strictly necessary for this paper. Section 3 contains the precise definitions and statements of our main results as well as their proofs. Finally, we add a short appendix to show that lca groups always satisfy property A’. This explains why unlike in the case of Špakula and Willett [23] no additional geometrical assumption is necessary.
@@
for $x,y\in\Xi$, cf. [1, Theorem 3.3] for the relevant uniqueness theorem and [10, Section 2] for more explanations on this. Since we assumed $m(e,e)=1$, it is not difficult to verify that $U_{x}^{\ast}=\overline{m(x,x^{-1})}U_{x^{-1}}$. Moreover, the cocycle condition and $m(e,e)=1$ imply $m(e,x)=1$ for all $x\in\Xi$.
@@
    The assumption that $m(x,y)=m(x^{-1},y^{-1})$ for all $x,y\in\Xi$ ensures that there exists a unitary operator $R\in\mathcal{U}(\mathcal{H})$ satisfying $RU_{x}=U_{x^{-1}}R$. $R$ can be chosen to be self-adjoint (cf. [10, Section 2]).
@@
    Note that the set $\mathcal{H}_{1}$ defined in the assumption is automatically dense in $\mathcal{H}$ and even contains a dense subspace of $\mathcal{H}$. This can be proven as for standard unitary representations, cf. [4, Proposition 14.5.1]. Further, it is itself indeed a subspace, as we will see later.
@@
We are now going to describe the basics of coorbit theory in the setting of our projective representations. We do not want to hide that our discussion of coorbit theory is based on the very nice survey [2], and most of the results between Lemma 2.2 and Proposition 2.16 are based on the analogous results in that survey without giving explicit reference at each occasion. Nevertheless, since, on the one hand, [2] does not deal with projective unitary representations and considers only “standard” unitary representations (in this case, many of the occurring formulas are much nicer), as well as the fact that [2] always carries the Duflo-Moore operator along (which is not necessary here, as we are working on unimodular groups), we decided to give full proofs and details to keep our presentation self-contained for the reader’s convenience.
@@
The previous lemma is just a reformulation of Godement’s orthogonality relations, cf. [10, Theorem 2.4]. We set $\|f\|_{2,\varphi_{0}}:=\|\mathcal{W}_{\varphi_{0}}(f)\|_{L^{2}(\Xi)}$. Furthermore, by the continuity properties of the representation, each element of $\mathcal{W}_{\varphi_{0}}(\mathcal{H})$ is a bounded and continuous function, which means that we can evaluate the functions at specific points.
@@
If $\phi\in\operatorname{Co}_{\infty}(U)$, then $\big{[}\mathcal{W}_{\varphi_{0}}(U_{x}\phi)\big{]}(y)=m(x,x^{-1})\overline{m(x^{-1},y)}\big{[}\mathcal{W}_{\varphi_{0}}(\phi)\big{]}(x^{-1}y)$.
@@
    Since $M^{2}(G)=L^{2}(G)$, multiplying by characteristic functions $\mathbbm{1}_{K}$ and $\mathbbm{1}_{G\setminus HK}$ is well-defined in $M^{2}(G)$ and therefore (3.1) is equivalent to $M_{\mathbbm{1}_{G\setminus HK}}AM_{\mathbbm{1}_{H}}=0$ in case $p=2$. In this form, Definition 3.2 is a metric-free version of [14, Definition 3.2]. More precisely, if we additionally assume that $G$ is second countable, then it is properly metrizable (in the sense that every closed ball is compact) and our definition of band operators becomes equivalent to the usual definition (e.g. [14, Definition 3.2] or [21, Definition 2.1.5]). The compact set $K$ can be chosen as a ball of a certain radius centered at $e$ and the smallest possible radius is then the band width of $A$ in the sense of [14, Definition 3.2]. For instance, if $G=\mathbb{Z}$ and $A\in\operatorname{BO}_{K}(\mathbb{Z})$, this means that only the diagonals $k\in K$ in the matrix representation of $A$ can be non-zero, where $k=0$ represents the main diagonal and so on.
@@
using that $M^{q}(G)^{\ast}\cong M^{p}(G)$ (respectively, $M^{1}(G)\cong(M^{\infty,0}(G))^{\ast}$; see Lemma 2.17), we see that $\mathcal{L}(M^{p}(G))$ is always a dual space. Hence, $\mu\ast A$ is always defined as a Gelfand integral, cf. [10, Section 3]. A key property that we will make use of is the fact that $\|\mu\ast A\|\leq\|\mu\|\cdot\|A\|$, where $\|\mu\|$ refers to the total variation norm of $\mu$.
@@
Hence, we can consider $A$ as a bounded linear operator from $S_{0}(G)$ to $S_{0}^{\prime}(G)$. By the outer kernel theorem for $S_{0}(G)$ (cf. [16, Theorem 9.3]), every bounded linear operator from $S_{0}(G)$ to $S_{0}^{\prime}(G)$ is represented by an integral kernel $k_{A}\in S_{0}^{\prime}(G\times G)$, given by the relation:
@@
For the inclusion $\supseteq$” let $K\subseteq G$ be compact and $A\in\operatorname{BO}_{K}^{p}(G)$. Let $g\in L^{1}(\widehat{G})$ be such that $\widehat{g}=1$ on $K$ (this always exists, cf. [22, Prop. 5.1.9]). Then, by Lemma 3.8, the kernel of $A$ satisfies
@@
Now, imitating the proof of [12, Lemma 3.8], we obtain the following result. Here, we write $\partial\Xi:=\mathcal{M}(\operatorname{BUC}(G\times\widehat{G}))\setminus(G\times\widehat{G})$.
@@
Here, we use $\alpha_{x}(B)$ and $\alpha_{\xi}(B)$ as a shorthand for the strong limits $\lim\limits_{x_{\gamma}\to x}\alpha_{(x_{\gamma},1)}(B)$ and $\lim\limits_{\xi_{\gamma}\to\xi}\alpha_{(e,\xi_{\gamma})}(B)$, respectively. Moreover, by $\partial G$ we technically mean the Stone–Čech boundary. However, since $G$ is discrete in this instance, $\partial G$ is isomorphic to $\mathcal{M}(\operatorname{BUC}(G))\setminus G$. In fact, for this corollary specifically, any compactification of $G$ can be used. Indeed, one might as well reformulate “$\alpha_{x}(B)=0$ for every $x\in\partial G$” to “$\alpha_{x_{\gamma}}(B)\to 0$ for any net $(x_{\gamma})_{\gamma\in\Gamma}$ that has no convergent subnet in $G$” and not use any compactification at all. For a brief discussion about different compactifications we refer to [14, Section 5].
@@ Appendix A Lca groups and Property A’
The following is a metric-free version of [14, Definition 2.3].
@@
As every lca group is amenable by [20, Proposition 0.15], [20, Theorem 4.10] implies that for every $\varepsilon>0$ and every compact subset $H\subseteq G$ there is a compact subset $K\subseteq G$ with $\lambda(K)>0$ such that
@@
Let $U\subseteq G$ be an open, relatively compact neighbourhood of $e$. Then $\left\{gU:g\in G\right\}$ is an open cover of $G$. By [15, Chapter II, Theorem 8.13], every lca group is paracompact. Therefore $\left\{gU:g\in G\right\}$ has a locally finite, open refinement $\left\{U_{j}:j\in\mathcal{J}\right\}$ for some index set $\mathcal{J}$. That is, for every compact set $K^{\prime}\subseteq G$ the set
@@
By [16, Lemma 4.3(iii)], we know that $\rho_{j}\in S_{0}(G)$. It is also clear that each $\rho_{j}$ has compact support. More precisely, if $g^{-1}h\notin K$ for all $h\in U_{j}$, then $\rho_{j}(g)=0$ and it follows that the support of $\rho_{j}$ is contained in $U_{j}K^{-1}=g^{\prime}UK^{-1}$ for some $g^{\prime}\in G$. Moreover, each $\rho_{j}$ is continuous since

### v2

@@ Abstract
AMS subject classification: Primary: 47L80; Secondary: 47B90; 47B07; 22D10; 22D25; 43A70
Keywords: quantum harmonic analysis, locally compact abelian groups, operator algebras, coorbit theory, compactness characterization, limit operators
@@ Introduction
where $e$ is the neutral element of the group $\Xi$. For $G:=\mathbb{R}^{d}$ (and hence $\Xi=\mathbb{R}^{2d}$), $\mathcal{C}_{1}(\mathcal{H})$ has spectral properties similar to band-dominated operators, cf. [13, 15]. In fact, it is well known that in case $G=\mathbb{Z}^{N}$ (hence $\Xi=\mathbb{Z}^{N}\times\mathbb{T}^{N}$) the algebra of band-dominated operators actually agrees with $\mathcal{C}_{1}(\ell^{2}(\mathbb{Z}))$, cf. [25, Theorem 2.1.6].
@@
Indeed, what we are going to prove is a bit more detailed than the above statements, and the precise formulation needs some more technicalities that we only mention in later parts of the paper. To briefly summarize this, we prove the result not only for the Hilbert space $L^{2}(G)$, but also for operators on the modulation spaces $M^{p}(G)$. If the group $G$ is discrete, then $M^{p}(G)=\ell^{p}(G)$, putting the results in alignment with the classical results for band-dominated operators over sequence spaces, noting that $\operatorname{BDO}(M^{p}(\widehat{G}))=\mathcal{L}(M^{p}(\widehat{G}))$ and $\partial\Xi=\partial G\times\widehat{G}$ in this case. When the group is not discrete, then $M^{p}(G)\neq L^{p}(G)$ for $p\neq 2$. Further, we are actually able to classify the band-dominated operators (and not only their intersection with Fourier-BDO) in terms of continuity with respect to a certain group action, leading to the above theorem. Let us also mention that we do not assume that $G$ is metrizable. In fact, we do not even need first countable. This contrasts to previous work such as [18] and [27], where the metric is an essential part of the theory. Instead of a metric, we will use the uniform structure that the lca group provides. We note that this involves some technical issues when the group is not second countable. Nevertheless, all these problems can be overcome, mainly by making an appropriate change in the definition of what the space $L^{\infty}$ is. We shall not elaborate on the details of this here and instead refer to the discussions in [14, End of Section 3] or [11, Section 2.3].
The paper is organized as follows. In Section 2, we describe the basic setting of the present work, combined with an introduction to coorbit spaces, as suitable for the precise setting we are working with. Since we also want to make this section available as a reference for later work, we put some more details there than would be strictly necessary for this paper. Section 3 contains the precise definitions and statements of our main results as well as their proofs. We also provide some simple examples to indicate how our results can be applied. Finally, we add a short appendix to show that lca groups always satisfy property A’. This explains why unlike in the case of Špakula and Willett [27] no additional geometrical assumption is necessary.
@@
for $x,y\in\Xi$, cf. [1, Theorem 3.3] for the relevant uniqueness theorem and [14, Section 2] for more explanations on this. Since we assumed $m(e,e)=1$, it is not difficult to verify that $U_{x}^{\ast}=\overline{m(x,x^{-1})}U_{x^{-1}}$. Moreover, the cocycle condition and $m(e,e)=1$ imply $m(e,x)=1$ for all $x\in\Xi$.
@@
    The assumption that $m(x,y)=m(x^{-1},y^{-1})$ for all $x,y\in\Xi$ ensures that there exists a unitary operator $R\in\mathcal{U}(\mathcal{H})$ satisfying $RU_{x}=U_{x^{-1}}R$. $R$ can be chosen to be self-adjoint (cf. [14, Section 2]).
@@
    Note that the set $\mathcal{H}_{1}$ defined in the assumption is automatically dense in $\mathcal{H}$ and even contains a dense subspace of $\mathcal{H}$. This can be proven as for standard unitary representations, cf. [5, Proposition 14.5.1]. Further, it is itself indeed a subspace, as we will see later.
@@
We are now going to describe the basics of coorbit theory in the setting of our projective representations. The theory of coorbit spaces has been developed in the papers [8, 9, 10] by H. G. Feichtinger and K. Gröchenig. A detailed discussion for the case of projective representations has been given in [3].
In the following, we are going to describe the essentials of the theory of coorbit theory that are necessary for the following sections of the paper. Since we do not assume the reader to be familiar with this theory, we decided to formulate the results in precisely the setting we are working in (in contrast to the more general works cited above) and to give full proofs of all these results. We do not want to hide that our discussion of coorbit theory is based on the very nice survey [2], and most of the results between Lemma 2.2 and Proposition 2.16 are based on the analogous results in that survey without giving explicit reference at each occasion.
@@
The previous lemma is just a reformulation of Godement’s orthogonality relations, cf. [14, Theorem 2.4]. We set $\|f\|_{2,\varphi_{0}}:=\|\mathcal{W}_{\varphi_{0}}(f)\|_{L^{2}(\Xi)}$. Furthermore, by the continuity properties of the representation, each element of $\mathcal{W}_{\varphi_{0}}(\mathcal{H})$ is a bounded and continuous function, which means that we can evaluate the functions at specific points.
@@
For $\phi\in\operatorname{Co}_{\infty}(U)$ we have
$\big{[}\mathcal{W}_{\varphi_{0}}(U_{x}\phi)\big{]}(y)=m(x,x^{-1})\overline{m(x^{-1},y)}\big{[}\mathcal{W}_{\varphi_{0}}(\phi)\big{]}(x^{-1}y).$
@@
    Since $M^{2}(G)=L^{2}(G)$, multiplying by characteristic functions $\mathbbm{1}_{K}$ and $\mathbbm{1}_{G\setminus HK}$ is well-defined in $M^{2}(G)$ and therefore (3.1) is equivalent to $M_{\mathbbm{1}_{G\setminus HK}}AM_{\mathbbm{1}_{H}}=0$ in case $p=2$. In this form, Definition 3.2 is a metric-free version of [18, Definition 3.2]. More precisely, if we additionally assume that $G$ is second countable, then it is properly metrizable (in the sense that every closed ball is compact) and our definition of band operators becomes equivalent to the usual definition (e.g. [18, Definition 3.2] or [25, Definition 2.1.5]). The compact set $K$ can be chosen as a ball of a certain radius centered at $e$ and the smallest possible radius is then the band width of $A$ in the sense of [18, Definition 3.2]. For instance, if $G=\mathbb{Z}$ and $A\in\operatorname{BO}_{K}(\mathbb{Z})$, this means that only the diagonals $k\in K$ in the matrix representation of $A$ can be non-zero, where $k=0$ represents the main diagonal and so on.
@@
using that $M^{q}(G)^{\ast}\cong M^{p}(G)$ (respectively, $M^{1}(G)\cong(M^{\infty,0}(G))^{\ast}$; see Lemma 2.17), one can see that $\mathcal{L}(M^{p}(G))$ is always a dual space. Hence, $\mu\ast A$ is always defined as a Gelfand integral, cf. [14, Section 3]. A key property that we will make use of is the fact that $\|\mu\ast A\|\leq\|\mu\|\cdot\|A\|$, where $\|\mu\|$ refers to the total variation norm of $\mu$.
@@
Hence, we can consider $A$ as a bounded linear operator from $S_{0}(G)$ to $S_{0}^{\prime}(G)$. By the outer kernel theorem for $S_{0}(G)$ (cf. [20, Theorem 9.3]), every bounded linear operator from $S_{0}(G)$ to $S_{0}^{\prime}(G)$ is represented by an integral kernel $k_{A}\in S_{0}^{\prime}(G\times G)$, given by the relation:
@@
For the inclusion $\supseteq$” let $K\subseteq G$ be compact and $A\in\operatorname{BO}_{K}^{p}(G)$. Let $g\in L^{1}(\widehat{G})$ be such that $\widehat{g}=1$ on $K$ (this always exists, cf. [26, Prop. 5.1.9]). Then, by Lemma 3.8, the kernel of $A$ satisfies
@@
Now, imitating the proof of [16, Lemma 3.8], we obtain the following result. Here, we write $\partial\Xi:=\mathcal{M}(\operatorname{BUC}(G\times\widehat{G}))\setminus(G\times\widehat{G})$.
@@
Here, we use $\alpha_{x}(B)$ and $\alpha_{\xi}(B)$ as a shorthand for the strong limits $\lim\limits_{x_{\gamma}\to x}\alpha_{(x_{\gamma},1)}(B)$ and $\lim\limits_{\xi_{\gamma}\to\xi}\alpha_{(e,\xi_{\gamma})}(B)$, respectively. Moreover, by $\partial G$ we technically mean the Stone–Čech boundary. However, since $G$ is discrete in this instance, $\partial G$ is isomorphic to $\mathcal{M}(\operatorname{BUC}(G))\setminus G$. In fact, for this corollary specifically, any compactification of $G$ can be used. Indeed, one might as well reformulate “$\alpha_{x}(B)=0$ for every $x\in\partial G$” to “$\alpha_{x_{\gamma}}(B)\to 0$ for any net $(x_{\gamma})_{\gamma\in\Gamma}$ that has no convergent subnet in $G$” and not use any compactification at all. For a brief discussion about different compactifications we refer to [18, Section 5].
@@
Let us now consider some simple examples of operators on $L^{2}(G)$.
Example 3.16.
  • (a)
    As mentioned in the introduction, it appears that the limit operator method does not work as expected for multiplication operators on $\mathbb{R}$. We want to revisit this example to see what is going wrong. Let $G$ be an arbitrary lca group and let us write $\Xi:=G\times\widehat{G}$ as before. Consider $f\in C_{0}(G)$ such that $\widehat{f}\in L^{1}(\widehat{G})$. Then clearly the multiplication operator $M_{f}$ is in $\operatorname{BDO}^{2}(G)$. Moreover,
    $C_{\widehat{f}}g(\nu):=\mathcal{F}M_{f}\mathcal{F}^{-1}g(\nu)=\int_{\widehat{G}}\widehat{f}(\mu^{-1}\nu)g(\mu)\,\mathrm{d}\mu$
    for all $g\in L^{2}(\widehat{G})$ and $\nu\in\widehat{G}$. As we assumed $\widehat{f}\in L^{1}(\widehat{G})$, Lemma 3.8 shows that $\mathcal{F}M_{f}\mathcal{F}^{-1}\in\operatorname{BDO}^{p}(\widehat{G})$. Now let us consider the limit operators of $M_{f}$. As expected,
    $\alpha_{(x,\nu)}(M_{f})=M_{f(x^{-1}\cdot)}$
    and so if we fix $\nu\in\widehat{G}$, then $\alpha_{(x,\nu)}(M_{f})\to 0$ as $(x,\nu)\to\partial\Xi$. However, if we fix $x\in G$, say $x=e$, then $\alpha_{(e,\nu)}(M_{f})$ does not converge to 00 as $(e,\nu)\to\partial\Xi$. Therefore Theorem 3.11 implies that $M_{f}$ is not compact unless $G$ is discrete (because in the latter case there is no net $(\nu_{\gamma})_{\gamma\in\Gamma}$ such that $(e,\nu_{\gamma})\to\partial\Xi$, see Corollary 3.12). Hence we get the expected result.
  • (b)
    Let $\varphi\in C_{0}(G)$ and $\psi\in L^{1}(G)$ such that $\widehat{\varphi}\in L^{1}(\widehat{G})$ and $\widehat{\psi}\in C_{0}(\widehat{G})$. Consider the product of the multiplication operator $M_{\varphi}$ and the convolution operator $C_{\psi}$:
    $A:=M_{\varphi}C_{\psi}.$
    Note that $A$ is an integral operator but not Hilbert–Schmidt in general. Then as in Example (a) one can easily see that both $M_{\varphi}$ and $C_{\psi}$ are in $\operatorname{BDO}^{p}(G)$ and in $\mathcal{F}^{-1}\operatorname{BDO}^{p}(\widehat{G})\mathcal{F}$. This of course implies $A\in\operatorname{BDO}^{p}(G)\cap(\mathcal{F}^{-1}\operatorname{BDO}^{p}(\widehat{G})\mathcal{F})$. Let $(x_{\gamma},\nu_{\gamma})_{\gamma\in\Gamma}$ be a net in $\Xi$ that converges to some $z\in\partial\Xi$. Since
    $\alpha_{(x,\nu)}(M_{\varphi})=M_{\varphi(x^{-1}\cdot)}\quad\text{and}\quad\alpha_{(x,\nu)}(C_{\psi})=C_{\psi\nu},$
    we either get $\alpha_{z}(M_{\varphi})=0$ if $(x_{\gamma})_{\gamma\in\Gamma}$ escapes every compact set or $\alpha_{z}(C_{\psi})=0$ if $(\nu_{\gamma})_{\gamma\in\Gamma}$ escapes every compact set. Either way, we have $\alpha_{z}(A)=0$ for all $z\in\partial\Xi$. Theorem 3.11 thus implies that $A$ is compact.
  • (c)
    More generally, we can consider integral operators $A\colon L^{2}(G)\to L^{2}(G)$ with suitably nice kernel $k_{A}\colon G\times G\to\mathbb{C}$. Then $\widetilde{A}:=\mathcal{F}A\mathcal{F}^{-1}\colon L^{2}(\widehat{G})\to L^{2}(\widehat{G})$ is also an integral operator with kernel
    $k_{\widetilde{A}}(\nu,\mu)=\int_{G}\int_{G}k_{A}(x,y)\overline{\nu(x)}\mu(y)\,\mathrm{d}x\,\mathrm{d}y.$
    Hence, for instance, if there are $g\in L^{1}(G)$ and $h\in L^{1}(\widehat{G})$ such that $\left\lvert k_{A}(x,y)\right\rvert\leq g(y^{-1}x)$ and $\left\lvert k_{\widetilde{A}}(\nu,\mu)\right\rvert\leq h(\mu^{-1}\nu)$, then $A\in\operatorname{BDO}^{p}(G)\cap(\mathcal{F}^{-1}\operatorname{BDO}^{p}(\widehat{G})\mathcal{F})$. Moreover,
    $\alpha_{(x,\nu)}(A)f(y)=\int_{G}k_{A}(x^{-1}y,x^{-1}z)\nu(z^{-1}y)f(z)\,\mathrm{d}z.$
    The compactness condition $\alpha_{(x,\nu)}(A)\to 0$ as $(x,\nu)\to\partial\Xi$ thus translates to a condition on the kernel of $A$. This very much resembles the matrix case discussed in Theorem 3.15.
@@ Appendix A Lca groups and Property A’
The following is a metric-free version of [18, Definition 2.3].
@@
As every lca group is amenable by [24, Proposition 0.15], [24, Theorem 4.10] implies that for every $\varepsilon>0$ and every compact subset $H\subseteq G$ there is a compact subset $K\subseteq G$ with $\lambda(K)>0$ such that
@@
Let $U\subseteq G$ be an open, relatively compact neighbourhood of $e$. Then $\left\{gU:g\in G\right\}$ is an open cover of $G$. By [19, Chapter II, Theorem 8.13], every lca group is paracompact. Therefore $\left\{gU:g\in G\right\}$ has a locally finite, open refinement $\left\{U_{j}:j\in\mathcal{J}\right\}$ for some index set $\mathcal{J}$. That is, for every compact set $K^{\prime}\subseteq G$ the set
@@
By [20, Lemma 4.3(iii)], we know that $\rho_{j}\in S_{0}(G)$. It is also clear that each $\rho_{j}$ has compact support. More precisely, if $g^{-1}h\notin K$ for all $h\in U_{j}$, then $\rho_{j}(g)=0$ and it follows that the support of $\rho_{j}$ is contained in $U_{j}K^{-1}=g^{\prime}UK^{-1}$ for some $g^{\prime}\in G$. Moreover, each $\rho_{j}$ is continuous since